OpenStudy (anonymous):

What is the sum of the geometric sequence 1, 3, 9, … if there are 14 terms

OpenStudy (campbell_st):

well the 1st term is a = 1 the common ratio is r = 3 and the number of terms n = 19 so you need to substitute them into the formula \[S_{n} = \frac{a(r^n - 1)}{r - 1} \] to get the answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

plz explain in depth

OpenStudy (campbell_st):

do you understand common ratio...

OpenStudy (anonymous):

cause i ain't getting the answer and yeah

OpenStudy (campbell_st):

ok... so here is the substitution \[S_{19} = \frac{1 \times ( 3^{19} - 1)}{ 3 - 1}\] now evaluate

OpenStudy (anonymous):

were do you get 19 from ???

OpenStudy (campbell_st):

oops that should be 14 \[S_{14} = \frac{ 1 \times (3^{14} - 1)}{3 -1}\]

OpenStudy (anonymous):

yeah can you give me the answer i am trying it and doesn't seem right

OpenStudy (campbell_st):

here's a calculator that may help http://web2.0calc.com/

OpenStudy (anonymous):

i have a graphing cal and a scientific still formual isn't comeing out right for my multiple choice

OpenStudy (anonymous):

1,285,956 2,391,484 3,556,228 4,287,382

OpenStudy (campbell_st):

all you need to do on you calculator is enterit as its written below \[1 \times (3^{14} -1)\div(3 -1)\] then press equals..

OpenStudy (anonymous):

thanks for being little more clear

OpenStudy (anonymous):

Ty!! :) This was exaplained really clear, just simplfy 1*(3^14-1)/(3-1) :D